OpenStudy (anonymous):

kris is 1.5 meters tall and casts a shadow 4 meters long. at the same time, a statue casts a shadow 12 meters long. what is the height of the statue? one

OpenStudy (anonymous):

\[\frac{1.5}{4}=\frac{x}{12}\iff x=\frac{12\times 1.5}{4}\]